OpenStudy (anonymous):

MEDAL + FAN AWARDED FOR BEST ANSWER! Find a situation in your everyday life that you can use unit rates to compare. Explain the situation, include calculations to show the unit rates, and determine which is the better rate. Here are some examples, but you may think of your own example: Number of channels offered by different cable companies compared to the monthly cost, Cell phone rates, minutes spent on your science and history homework compared to the number of problems, the speed you complete a race, batting average or comparison shopping

OpenStudy (anonymous):

@TylerD @timeflybyegirl

OpenStudy (anonymous):

nvm i got one

OpenStudy (anonymous):

is this good?

OpenStudy (anonymous):

1 hour to complete 4 math exams 60/4 = 15 that would be 15 minutes per exam

OpenStudy (tylerd):

you have to compare 2 things that would be one

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

ok here

OpenStudy (anonymous):

1 hour to complete 4 math exams 1 hour to complete 2 science assignments Math __________________________ 60/4 = 15 that would be 15 minutes per exam ___________________________ Science ____________________________ 60 / 2 = 30 which would be 30 minutes per assessment/assignment. _____________________________________________________________

OpenStudy (anonymous):

is that good @TylerD

OpenStudy (anonymous):

And thats about how long it takes me

OpenStudy (tylerd):

could work if you can come up with some kind of conclusion

OpenStudy (anonymous):

ok

OpenStudy (tylerd):

if you wanted to do something a bit more challenging you could google cell phone plans, and explain which one is cheaper. or something like that.

OpenStudy (anonymous):

nah i think ill be fine

OpenStudy (anonymous):

my teacher is nice :P

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

An animal shelter is deciding between two brands of cat food. Two tubs of cat food are shown. The first one is Happy Cat. It costs forty-eight dollars for fifteen pounds. The second one is Cat's Meow. It costs fifty dollars for sixteen pounds. The shelter is happy with both brands of cat food, so they will make their decision based on price. Which brand of cat food is the better buy? Explain how you know, showing all your work.